# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

### v2

@@ 6.2.3. Racial disparities in a pandemic
6.3. Non-COVID CDC guidelines
7. Study III: Tweeting through the pandemic
In Study II (§6), we found that periods of policy change introduced more uncertainty to the public. For example, clusters 2, 4, 7 (focusing on the lack of clarity in CDC guidelines), 13, 14, 15, and 16 all focused on policy change, like removing the need to mask after receiving the second COVID vaccine shot, a policy that was then changed with the Omicron variant wave (cluster 4). Given that it is important to understand flashpoints in crisis management (see §2.3.3), we used temporal models including change-points and survival analysis (detailed in §4.4)) to analyze: (1) main flashpoints over more than two-years of the pandemic (§7.1 and §7.2); and (2) behavior that increased the chance they would remain engaged in the CDC public (§7.3).
Summary of findings:
In §7.1, we describe three COVID crisis phases based on change-point that tracked early discussions about the seriousness of COVID, the introduction of mitigation guilines, and changing them with new variants. We then showed how the pattern of rich media use coincided with changes in CDC messaging about mitigation (e.g., the return to masking after vaccinations due to new variants). Finally, we showed how the use of rich media, being verified, sharing low-credibility sources, and being retweeted all contributed to long-term engagement on the CDC public.
7.1. Evolving discourse on CDC-related publics: three phases of the COVID crisis
In this section, we report our findings on pandemic phases using change-points as detailed in section 4.4.1. In our findings, we show that there there three main phases: (1) hoaxes and early pandemic effects; (2) lock-downs and masks; and (3) post-vaccine variants and evolving guidelines.
7.1.1. Beginning Phase: Hoaxes and early pandemic effects
Early on in the pandemic, discussions about the race disparities in healthcare (see Figure 7) and communication about the dangers of vaping (Figure 6 both dropped significantly and remained low throughout the period of the study. This shows that in the early days of 2020, COVID is becoming more of an emergency with LLR (see 4.4 for details) values showing that Topics 62 (social distancing guidelines), 59 (parody of CDC guidelines), 57 (facial hair and mask guidelines), and 34 (news reflecting the role of the CDC in managing the pandemic) become more prevalent after the first change-point.
While the racial disparities theme showed a change-point trending downwards early in the pandemic, we have identified a cluster focused on race challenges in the pandemic, described in more detail in §6.2.3. To determine if there were any change-points specific to the racial disparity cluster, we repeated the PELT analysis, specifically for the Cluster 1 subset. This is presented in Figure 7. As you can see, there is a continuous focus on Topic 40: COVID racial disparities throughout most the duration of this analysis (note that the duration for the cluster is not the same as the rest of the dataset because it covers a subset of 121 users). This shows that flashpoints detected on the community level could be different than those that represent a specific set of stakeholders forming an interest public of their own - in this case, the healthcare equity public.
The next change-point is for Topic 1: Using COVID data to counter the idea that it is a hoax. We found that there are two change points, one on the 21st of January 2020 and the other is on 31st 2020. Both these change-points show a downward and significant trend suggesting that this topic peaked between two dates. Topic 48 (Hydroxychloroquine and Ivermectin controversy) was one of the main topics associated with the first change-point. As the name suggests, this topic focused on the veracity (or lack thereof), of the use of these two drugs to treat COVID. While some Tweets were sharing information showing their lack of efficacy, less traditional links were also shared about the need to use these medications that are available on the market.
Both change points for Topic 1 increased the discussion of Topic 62 (social distancing guidelines). Topics 59 (parody of CDC guidelines), 57 (facial hair and mask guidelines), and 34 (news reflecting the role of the CDC in managing the pandemic) are associated with both change-points. These topics focus on the guidelines shared by the CDC as the effects of the COVID emergency become clearer. While some Tweets focused their framing on taking COVID more seriously given the seriousness of COVID,
Finally, one of the change-points in the Testing Theme (see Figure 8) focused on the lack of available COVID tests and the low accuracy of the available test (see details in §6.2.1).
Figure 6. This figure shows change-points for Topic 1 (using data to counter idea of hoax), and Topic 45 (vaping dangers), the mask them. As evident in the figure, all three have significant change-points early in the pandemic. The mask theme has a second change-point discussed in the second phase. Green dots show significant trends and blue arrows show trend directions.
7.1.2. Middle Phase: Lock-downs and masks
The introduction of lock-downs and the call for masking guidelines represent middle part of the COVID crisis. The discourse here started with a discussion of the lack of direction from the CDC (parodies of CDC guidelines).
After the second change-point by late March for Topic 1, the discourse shifted more towards Topic 30: in-person gathering guidelines. This shows how after the first lock-down order was issued in the US on March 15th of 2020, the focus became more on managing the risks associated with COVID, rather than discussions around whether COVID is a hoax, or making light of changing CDC guidelines. However, some posts took a different approach framing the CDC’s guidance as being more draconian. One of the areas of discussion was the suggestion of masking outdoors or double masking (e.g., wearing two masks). Others criticized what they saw as the constantly changing guidelines coming out of the CDC. This is especially true for masking.
Discussions about masking peaked between March 31st 2020 and June 14th 2021 (see Figure 6). Discourse around the first change-point early in the pandemic is focused on guidelines on wearing masks and the CDC’s role in managing the pandemic (Topics 30, 57, 62, 34, and 59). After the first change-point, many Tweets focused on the fact that the CDC changed its position on masking from the early days of the crisis when the guidance was that only healthcare workers and anyone showing symptoms should mask while later on, masking was suggested for all in-doors activities. One example comparing earlier guidelines is presented in Figure 19. We discuss the second masking change-point in §7.1.3.
Figure 7. While there is an early change-point showing that Race discourse decreased early on in the pandemic, COVID Racial Disparities (Topic 40) is heavily discussed in Cluster 1, showing that change-points vary across different publics.
7.1.3. End Phase: Post-vaccine variants and evolving guidelines
Towards the tail of the period covered in this study, we found that change-points are detected in connection to: (1) vaccination; and (2) new variants. The first change-point is about Topic 0: Voices for and against vaccinations. The significant change-point was detected on December 26th 2021. This period of time coincided with the hight of the Omicron peak (as shown in Figure 9). Using the LLR signatures for discourse post change-point, we found that the two top topics signifying post-change-point discourse are Topics 47 (Omicron variant changing masking guidelines) and 31 (Delta variant changing masking guidelines). Both these topics focused on discussing if the new variants reduce the efficacy of the vaccine, and by reducing the efficacy of the vaccine, require the re-introduction of masking guidelines. Mitigation guidelines the focus of Topics 30 (in-person gatherings guidelines), 38 (effects of closures and re-openings on the economy), and 62 (social distancing guidelines), all of which were in flux because new variants reduced the efficacy of the COVID vaccines.
Similarly, discourse after the Mask Theme’s second change-point on June 14th (see Figure 6) focused on the evolving guidelines due to new variants (Topic 47: Omicron variant changing masking guidelines, Topic 31: Delta variant changing masking guidelines, Topic 62: social distancing guidelines). The CDC’s messaging about the changes associated with the new variants as they reduce the effectiveness of the vaccines are seen by many as confusing. They point to images like Figures 17 and18 where, while the CDC is reflecting how the new variants are less effective with the vaccine, and thus might require wearing masks especially in-doors, the claim is still that the vaccine is still effective against hospitalizations (see Figure 15).
A similar pattern is true for the School Management Post-COVID theme. The early change-point focused on discussions around social distancing and masking as the first school-year was being planned after the school closures early in the pandemic. The second change-point focused on the evolving guidelines and challenges associated with new variants. An example of this is shown in Figure 13 where the CDC image shows how the Delta variant is changing CDC guidelines for schools specifically because the vaccine is not as effective as it is to the wild-type virus.
Figure 8. This figure shows change-points in the medium phase and later phases, showing Topic 0 (voices for and against vaccination), COVID Test Theme, and the School Management Theme. Green dots show significant trends and blue arrows show trend directions.
Figure 9. COVID cases in Delta and Omicron waves
7.2. Change-points in CDC Tweet frequency, frequency of rich media content, and overall sentiment
When looking at the use of rich media in Tweets, we found two change-points, one early in pandemic on the 25th of February 2020, and then again on December 4th 2020. Both these change-points are significant with an upward trend. the LLR showing the top topics after the first change-point shows that pictures were first being used to relay detailed information about mitigation guidelines (Topic 62: social distancing guidelines, Topic 57: Facial hair and masking guidelines, Topic 48: Hydroxychloroquine and Ivermectin controversy).
The second change-point with rich media increases focus on Topics 4, 31, and 47 (Delta and Omicron variants changing masking guidelines) as well as evolving masking guidelines with vaccination where the CDC and other public health accounts were calling on the public to mask again since the new variants render the vaccine less effective. These change-points are in-line with the findings presented in §7.1. The change-points associated with “receipts,” usually screenshots of earlier Tweets from the CDC that might have provided different guidelines.
As the PELT timeseries in the middle of Figure 10 shows, Tweets by the CDC dropped to a minimum between February 25th 2020 and late May 2020. The trend after the change-points is positive, indicating that Tweets from the CDC were on the rise. However, between the two change-points, the topical focus was on criticizing the CDC. Specifically, making a parody of CDC masking guidelines (Topic 59) and the lack of their engagement on issues like the Hydroxychloroquine and Ivermectin controversy (Topic 48). This shows that when there was a vacuum in messaging from the CDC about the medical emergency, other members of the CDC public filled that message void by critiquing the lack of engagement.
The mean sentiment in the CDC public was positive until a change-point on November 1st 2020. The significant trend of after the change-point is negative. Topic 29 (Political/legislative COVID response) is the top LLR value after this change-point showing that there is a higher focus on the politics of COVID - from the mandatory introduction of vaccines, to reducing the number of lock-downs, to discussing whether government should be introducing a support package. One of the reasons for this might be the continued cycle of lock-downs and re-openings by this point. Specifically, there was significant push-back against in-person gathering guidelines for thanksgiving (see Figure 21). While this came later, Topic 47 (Omicron variant changing masking guidelines) also had a high LLR after this change-point, showing that reverting back to more stringent mitigation guidelines lowered the average sentiment in the CDC public.
Figure 10. This figure showing change-points for rich media, Tweets by the CDC, and overall Vader sentiment. Green dots show significant trends and blue arrows show trend directions
7.3. User commitment over time
In this section, we use the hazard ratio of the variables reported in Table 10 to determine the likelihood of their continued engagement in the CDC public. Any hazard ratio less than 1 increases the chance of their continued presence. The method is detailed in §4.4.2.
Being a verified user increases the likelihood of continued engagement (by 18%). Using rich media at a rate that is 1 standard deviation from average picture use is the most predictive of continuing engagement at 36% higher likelihood. The retweet variable shows that the those users who get retweets that are one standard deviation higher than average were 7% more likely to continue engaging in this discourse. Those users who use low-credibility sources of information at a rate one standard deviation higher than average, is 6% more likely to continue their engagement.
Interestingly, the higher the like count, the lower the likelihood of continual engagement (at 8%). Engaging in any of the main themes at a level 1 standard deviation higher than average reduce the likelihood of continued engagement (except for School and Vaccine themes which did not have any significant effects). When engaging in more positive sentiment at a value 1 standard deviation higher than average, the lower the likelihood of continued engagement in the community (6%).
These findings suggest that being a verified user, and using pictures in their Tweets, and being retweeted increases the chances of their continued engagement. While at a lower rates, the higher the use of low-quality information sources, and negative sentiment in their Tweets also increases the likelihood of continued engagement. On the other hand, focusing on any of the themes reduces the likelihood of continued engagement.
Table 10. Survival analysis predicting how long members continue to Tweet in the CDC public sphere. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05.
Covariate Type Hazard Ratio p
Retweet Count Platform signal 0.93 *
Low-credibility Behavior 0.94 ****
Rich Media Platform signal 0.64 ****
Verified Platform signal 0.82 ****
Like Count Platform signal 1.08 *
Masks Theme 1.03 ****
Mitigation Theme 1.02 ****
Non-COVID Theme 1.03 ****
Politics Theme 1.02 ****
Test Theme 1.02 ****
VADER Sentiment 1.06 ****

### v3

@@ 5.4.3. Racial disparities in a pandemic
Table 4. Cluster metadata including user counts, verified status, top biography descriptors (doc2vec and LLR), and E-I index. These indicators reflect ideological cohesion and engagement diversity within each cluster (part II).
# Cluster Name Users Top desc doc2vec Top desc LLR E-I Index
8 Obesity and COVID 87(2) Kamala, Streamer, hcaker, #research, conservative, stupid obesity, ASMBS, nutrition, dietitian, lbs, biostatician, diabetes, investigating, sugar -1.00
9 Gun violence 76(4) MAGA, vaccinated, Montanan, #Riseup, #GunReformNow, #Liberal securitywarriors, gunlawssavelives, endoflifeoptions, stopliars, votingrightsforall, trauma, momsdemand, 2nd -1.00
10 Food-borne diseases outbreak 278(11) Meat, Pastry, Trash bidenapocalypse, material, consumers
11 Vaccinating pregnant women 70(12) cweight, news10, childbirth, alabamastateu, newberrylibrary, iumedschool, mayocliniccv, umnmedschool, mildly @sloan_kettering, @EmoryUniversity, @News_8, #HealthEquity, #VaccinesSaveLives, Neuroscientist -1.00
12 School sports & community events 7760(584) accredited, supporter reporter, anchor, story, fox4kc, hiv, endorsements, uclaathletics, school, community, pediatric, comorbidity -0.24
13 Air travel guideline changes 209(23) Air Line Pilots Association, vacation, cbsaustin, thedailybeast @News_8, @NY1, Nuggets, Toronto, Correspondent, @sloan_kettering, @UCF, @HarvardMed, Epidemiology -1.00
15 School reopening dynamics 997(53) daughter, fall, accredited, childhood, counselor, curriculum, practitioner, athletic, vascular teacher, schools, educator, learner, education, superintendent, principal, mom, district, business owner, association -1.00
16 Quarantine dynamics 134(8) feet, tired, meaning rights, vol_baseball, vamosgalaxy, annmediasports, childhoodcancerawareness, oxy_football -1.00
5.5. Change Over Time: Comparing Marginalized and Broader Publics
The two PELT-detected time series reveal distinct temporal patterns in race-related discourse across the Twitter public (top) and Cluster 1 (bottom), which centers on COVID-related racial disparities. While race discourse in the broader public shows a sharp early decline, Cluster 1 maintains focused engagement with racial issues throughout the pandemic, as detailed in §5.4.3. This shows that flashpoints detected on the community level could be different than those that represent a specific set of stakeholders forming an interest public of their own - in this case, a healthcare equity public.
In the top panel, the time series shows a prominent early concentration of race-themed content, with a statistically significant changepoint detected on February 25, 2020, marking a sharp and sustained decline in discourse. The period before this changepoint (highlighted in blue) is characterized by high variability and frequent spikes, while the period after (highlighted in pink) reflects a flattened trend with minimal engagement—suggesting that race was a prominent topic early on but quickly diminished in focus within this cluster as the pandemic took hold. In contrast, the bottom panel reveals a more dynamic and episodic pattern, with multiple changepoints spanning from April 2019 to April 2021. This cluster displays intermittent but sharp increases in race-themed engagement—particularly around mid-2020—corresponding to periods of heightened racial awareness and public events, such as the George Floyd protests and subsequent public health debates. The alternating pink and blue segments indicate bursts of discourse rather than a consistent trend, suggesting that users in this cluster engaged with race themes in reaction to external events rather than as a sustained topic. We will focus on the discourse to provide examples for discourse in both.
”The first changepoint occurred between August 20, 2020 and October 24, 2020. The tweets in this breakpoint cluster around four intersecting themes that highlight how COVID-19 has deepened long-standing racial inequities in U.S. public health. Racial Disparities in COVID-19 Mortality and Hospitalization emphasize CDC data showing that Black, Hispanic, and Latinx individuals—especially children—face disproportionately high rates of COVID-19-related death and hospitalization. These tweets stress that such outcomes are not incidental but patterned and predictable. Intersection of Race, Age, and COVID-19 Impact highlights how younger adults, particularly those aged 25–44 from Hispanic and Latino communities, are also suffering disproportionately from the pandemic, challenging the perception that only older adults are vulnerable. Critique of Systemic Racism in Public Health brings attention to long-standing inequalities—such as drastically shorter life expectancies for Black individuals with Down syndrome compared to their white counterparts—and points to the broader discrediting of institutions as they become more racially diverse, exposing implicit racial bias. Finally, Calls for Political and Structural Accountability use CDC statistics as a foundation for civic and legislative demands. These tweets urge political leaders and voters to acknowledge and act on racial disparities, pushing for reforms that confront systemic injustice through policy change, taxation, and institutional responsibility.
The second changepoint was detected between April 2, 2021 and April 27, 2021. The discourse at this breakpoint reveals four interconnected themes centered on public health equity, racial justice, and institutional accountability. Data Justice & Visibility captures how several users challenge public health institutions—particularly the CDC—for aggregating health data in ways that obscure disparities. Tweets emphasize that grouping Native Hawaiian and Pacific Islander (NHPI) populations with all Asians, or generalizing Indigenous health statistics, erases the distinct health challenges these communities face. Public Health Advocacy for BIPOC Communities reflects coordinated efforts to promote COVID-19 vaccination within communities of color—especially among Latinos and Hispanics—using CDC data to highlight disproportionate infection rates. These messages are often accompanied by hashtags like #Vaccinate4All and endorsements from culturally aligned organizations such as @SaludAmerica and @NHMAmd. Institutional Racism & Accountability emerges as users share personal experiences and collective critiques of medical racism, calling on public health entities like the CDC and OSU Wexner Medical Center to acknowledge and address ongoing harm—particularly to Black women—and to end institutional gaslighting. Finally, Symbolic Actions vs. Structural Reform critiques the disjuncture between public-facing declarations—such as the CDC’s recognition of racism as a public health threat—and the lack of deep structural change across health and policy institutions. While symbolic gestures are noted, users call for more meaningful accountability and systemic reform.
Figure 5. Time-series of race-related discourse across the full Twitter public (top) and Cluster 1 (bottom). Early decline in race discourse contrasts with episodic but sustained engagement by Cluster 1. Detected change-points (red lines) show how this marginalized public responded dynamically to external events, unlike broader publics that deprioritized racial equity over time.